OpenStudy (anonymous):

what is the solution set of 10^x=2.89

OpenStudy (anonymous):

\[10^x=y \iff \log(y)=x\]

OpenStudy (anonymous):

do you know what the exact answer is? im stuck

OpenStudy (anonymous):

it is a calculator exercise. find \(\log(2.89)\) by pressing the appropriate buttons

OpenStudy (anonymous):

it says what is the solution in terms of natural logarithms? i got the 0.46 answer i just need the other solution set answer b4 that one

OpenStudy (zepp):

With the Change of base formula, we get \(\huge \frac{log2.89}{log10}\) base e \(\huge \frac{log_{e}2.89}{log_{e}10}=\frac{ln2.89}{ln10}\)

OpenStudy (anonymous):

thankyou

OpenStudy (precal):